# Multi-target Range, Doppler and Angle estimation in MIMO-FMCW Radar with Limited Measurements

### v1

@@ Appendix A Proof of Lemma 1
Consider the $(i_{1},i_{2})$-th element of $\mathbf{Q}_{B}$ as $[\mathbf{Q}_{B}]_{i_{1},i_{2}}$. Recall from Section IV that any column of matrix $\mathbf{B}$ is given by $\mathbf{b}(\rho_{i})=[\exp{(j\frac{4\pi T_{c}}{\lambda}\rho_{i}\zeta_{1})},\ldots,\exp{(j\frac{4\pi T_{c}}{\lambda}\rho_{i}\zeta_{P})}]^{T}$ such that
@@ Appendix B Proof of Theorem 1
Appendix B-A provides some preliminary results, including suitable bounds on $\Gamma_{B}$ and $\Gamma_{C}$. The bound on coherence $\mu$ is then derived in Appendix B-B.
@@ B-A Preliminaries
where $10<\vartheta<$ 1 and $i=2,3,\ldots,G_{D}$.
@@
Define $\zeta^{\prime}_{p}=\zeta_{p}-\frac{P_{max}-1}{2}$. Substituting this in (22), we obtain $\Gamma_{B}(u)=\exp{(ju\frac{P_{max}-1}{2})}\Gamma^{\prime}_{B}(u)$ where $\Gamma^{\prime}_{B}(u)=\frac{1}{P}\sum_{p=1}^{P}\exp{(ju\zeta^{\prime}_{p})}$. Now, since $\zeta_{p}\sim\mathcal{P}_{p}(\cdot)$, the random variable $\zeta^{\prime}_{p}$ has an even distribution because $\mathcal{P}_{p}(\cdot)$ is symmetric about $(P_{max}-1)/2$. Consequently, $\Gamma^{\prime}_{B}(u)$ is asymptotically jointly Gaussian distributed [56, Sec. II] as
$\begin{bmatrix}\textrm{Re}\;\Gamma^{\prime}_{B}\\ \textrm{Im}\;\Gamma^{\prime}_{B}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\textrm{Re}\;\Psi_{p^{\prime}}\\ \textrm{Im}\;\Psi_{p^{\prime}}\end{bmatrix},\begin{bmatrix}\sigma_{1}^{2}&0\\ 0&\sigma_{2}^{2}\end{bmatrix}\right),$
@@
$\begin{bmatrix}\textrm{Re}\;\Gamma^{\prime}_{B}\\ \textrm{Im}\;\Gamma^{\prime}_{B}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}0\\ 0\end{bmatrix},\begin{bmatrix}1/2P&0\\ 0&1/2P\end{bmatrix}\right),$
@@
Consider a MIMO radar with random SLA as described in Section II with $\alpha_{1:N_{T}}$ and $\beta_{1:N_{R}}$ drawn i.i.d. from even distributions $\mathcal{P}_{\alpha}$ and $\mathcal{P}_{\beta}$, respectively. If $\mathcal{P}_{\alpha}$, $\mathcal{P}_{\beta}$ and the uniform grid $\phi_{1:G_{\theta}}$ satisfy (26), then for $10<\vartheta<$ 1 and $i=2,3,\ldots,G_{\theta}$,
@@
The lemma is obtained trivially by generalizing the results of [31, Theorem 1] to the MIMO-FMCW radar array. ∎
@@ Appendix C Proof of Theorem 2
According to [52, Theorem 2.7], any $K$-sparse signal can be recovered if the measurement matrix $\mathbf{D}$ has RIP constant $\delta_{2K}<2/(3+\sqrt{7/4})\doteq\Lambda$. Hence, we need to bound $\delta_{2K}$ with a probability higher than $1-\epsilon$, or equivalently, $\mathbb{P}(\delta_{2K}>\Lambda)<\epsilon$ where $10<\epsilon<$ 1. However, RIP constant $\delta_{2K}$ and coherence $\mu$ of a matrix satisfy $\delta_{2K}\leq(2K-1)\mu$ [52]. Hence, $\mathbb{P}(\delta_{2K}>\Lambda)\leq\mathbb{P}(\mu>\Lambda/(2K-1))$.
@@
×$\times(1-2\vartheta_{0}\sqrt{N_{T}N_{R}}\;\mathcal{K}_{1}(2\vartheta_{0}\sqrt{N_{T}N_{R}}))^{G_{\theta}-1},$ (35)
@@
which can be solved using Lambert W function[58] as in the proof of [31, Theorem 2] to obtain (30). The constants $\kappa_{1}$ and $\kappa_{2}$ in (29)-(30) are obtained by substituting $\Lambda=2/(3+\sqrt{7/4})$.
For the case when $\alpha_{n}=\beta_{n}$, we have $\mathbb{P}(\delta_{2K}>\Lambda)<1-(1-e^{-\vartheta_{0}^{2}P})^{G_{D}-1}(1-e^{-N_{T}\vartheta_{0}})^{G_{\theta}-1}$ using (28) with $\vartheta_{0}=\Lambda/(2K-1)$. Again, using $(1-x)^{n}\approx 1-nx$ and $G_{D},G_{\theta}\gg 1$ and following similar steps as in the independent $\alpha_{1:N_{T}}$ and $\beta_{1:N_{R}}$ case, we obtain $G_{D}e^{-\vartheta_{0}^{2}P}+G_{\theta}e^{-N_{T}\vartheta_{0}}\leq\epsilon$, which is satisfied if $G_{D}e^{-\vartheta_{0}^{2}P}\leq\epsilon_{1}$ and $G_{\theta}e^{-N_{T}\vartheta_{0}}\leq\epsilon_{2}$ with $\epsilon_{1}+\epsilon_{2}=\epsilon$. Rearranging these inequalities, we obtain (29) and (31), respectively. Finally, the claim of the theorem then follows from [52, Theorem 2.7] considering the exact $K$-sparse signal case.
@@ Appendix D Proof of Theorem 3
If part: Using $\mathbf{D}_{N_{T}N_{R}(i-1)+j,:}=\mathbf{B}_{i,:}\otimes\mathbf{C}_{j,:}$, we obtain $\mathbb{E}[\mathbf{D}_{l,:}^{H}\mathbf{D}_{l,:}]=\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]\otimes\mathbb{E}[\mathbf{C}_{j,:}^{H}\mathbf{C}_{j,:}]$ for some $1\leq i\leq P$ and $1\leq j\leq N_{T}N_{R}$ because $\zeta_{1:P}$ is independent of $\alpha_{1:N_{T}}$ and $\beta_{1:N_{R}}$. We first consider $[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]$. Here, $\zeta_{1:P}$ are identically distributed such that $\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]$ does not depend on row index $i$ and $\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]=\frac{1}{P}\sum_{i=1}^{P}\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]$. But, by simple comparison of matrix elements, we can show that $\sum_{i=1}^{P}\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]=\mathbf{Q}_{B}$. Hence, $\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]=\widetilde{\mathbf{Q}}_{B}$ where $\widetilde{\mathbf{Q}}_{B}=(1/P)\mathbb{E}[\mathbf{Q}_{B}]$. As a consequence of Lemma 1, we only consider the first row of $\widetilde{\mathbf{Q}}_{B}$. Note that $(1,i)$-th element of $\mathbf{Q}_{B}$ is $\mathbf{B}_{:,1}^{H}\mathbf{B}_{:,i}=P\Gamma_{B}(u^{D}_{1,i})$ using (22). Hence, for $1\leq i\leq G_{D}$,
@@
where the second-last equality follows because $\zeta_{p}$ are i.i.d. for all $1\leq p\leq P$. Finally, if $\Psi_{p}(u^{D}_{1,i})=0$ for $i=2,3,\ldots,G_{D}$ according to (32), then $\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]=\widetilde{\mathbf{Q}}_{B}=\mathbf{I}_{G_{D}}$. Note that $[\widetilde{\mathbf{Q}}_{B}]_{1,1}=1$ because all main diagonal elements of $\mathbf{Q}_{B}$ is $P$ from Lemma 1. Further, following similar steps as in proof of [31, Theorem 3], we can trivially show that $\mathbb{E}[\mathbf{C}_{j,:}^{H}\mathbf{C}_{j,:}]=\mathbf{I}_{G_{\theta}}$ if $\Psi_{\xi}(u^{\theta}_{1,j})=0$ for $j=2,3,\ldots,G_{\theta}$ as in (32). Finally, $\mathbb{E}[\mathbf{D}_{l,:}^{H}\mathbf{D}_{l,:}]=\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]\otimes\mathbb{E}[\mathbf{C}_{j,:}^{H}\mathbf{C}_{j,:}]=\mathbf{I}_{G_{D}}\otimes\mathbf{I}_{G_{\theta}}=\mathbf{I}_{G_{D}G_{\theta}}$, i.e., if (32) holds, matrix $\mathbf{D}$ satisfies isotropy property.
Only if part: If $\Psi_{p}(u^{D}_{1,i})\neq 0$ for some $i=2,3,\ldots,G_{D}$, then $[\widetilde{\mathbf{Q}}_{B}]_{1,i}\neq 0$. Hence, $\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]\neq\mathbf{I}$ which implies $\mathbb{E}[\mathbf{D}_{l,:}^{H}\mathbf{D}_{l,:}]\neq\mathbf{I}$, i.e., matrix $\mathbf{D}$ does not satisfy isotropy property. Similarly, matrix $\mathbf{D}$ also does not satisfy isotropy property if $\Psi_{\xi}(u^{\theta}_{1,j})\neq 0$ for some $j=2,3,\ldots,G_{\theta}$. Therefore, (32) is a necessary condition.
@@ Appendix E Proof of Proposition 5
Note that $\alpha_{1:N_{T}},\beta_{1:N_{R}}\in[-1/2,1/2]$ when $A_{T}=A_{R}=A/2$ and the characteristic function of uniform distributions $\mathcal{P}_{\alpha}$ and $\mathcal{P}_{\beta}$ are
$\Psi_{\alpha}(u)=\Psi_{\beta}(u)=\frac{\sin(u/2)}{u/2}.$
From (26), we need $\Psi_{\alpha}(u^{\theta}_{i,j})=\Psi_{\alpha}(2u^{\theta}_{i,j})=0$ and $\Psi_{\beta}(u^{\theta}_{i,j})=\Psi_{\beta}(2u^{\theta}_{i,j})=0$ for $i\neq j$ which can be satisfied if $u^{\theta}_{i,j}/2=k\pi$ for some integer $k$. But, by definition, $\frac{u^{\theta}_{i,j}}{2}=\pi\frac{A}{2\lambda}(\sin{\phi_{j}}-\sin{\phi_{i}})$. Hence, condition (26) is satisfied if $\phi_{1:G_{\theta}}$ is a uniformly spaced grid of spacing $2\lambda/A$ in the $\sin{\theta}$ domain.

### v2

@@ VI Summary
While the proposed framework demonstrates promising performance in simulation, several practical challenges must be addressed for real-world deployment. Hardware imperfections, such as oscillator phase noise, ADC quantization errors, RF non-linearities, and mutual coupling between antennas, can distort the idealized measurement model. Additionally, environmental factors, including multipath propagation, target occlusion, and clutter, introduce further complications. Large aperture arrays may also encounter spatial wideband and near-field effects, leading to range-angle coupling and non-planar wavefronts, which complicate the estimation process and contribute to model mismatch, thereby degrading sparse recovery performance. To address these issues, future research will focus on developing calibration-aware and model-adaptive CS techniques, robust dictionary designs that account for practical distortions, and estimation methods resilient to noise and hardware impairments. We plan to experimentally validate the proposed framework using the integrated radar evaluation module platforms to assess its effectiveness under realistic hardware and propagation conditions.
@@ Appendix A Proof of Lemma 1
Consider the $(i_{1},i_{2})$-th element of $\mathbf{Q}_{B}$ as $\mathbf{Q}_{B}]_{i_{1},i_{2}}$. Recall from Section IV (of the main paper) that any column of matrix $\mathbf{B}$ is given by $\mathbf{b}(\rho_{i})=[\exp{(j\frac{4\pi T_{c}}{\lambda}\rho_{i}\zeta_{1})},\ldots,\exp{(j\frac{4\pi T_{c}}{\lambda}\rho_{i}\zeta_{P})}]^{T}$ such that
@@ Appendix B Proof of Theorem 1
In the following, Section B-A provides some preliminary results, including suitable bounds on $\Gamma_{B}$ and $\Gamma_{C}$. The bound on coherence $\mu$ is then derived in Section B-B.
@@ B-A Preliminaries
where $0<\vartheta<1$ and $i=2,3,\ldots,G_{D}$.
@@
Define $\zeta^{\prime}_{p}=\zeta_{p}-\frac{P_{max}-1}{2}$. Substituting this in (23), we obtain $\Gamma_{B}(u)=\exp{(ju\frac{P_{max}-1}{2})}\Gamma^{\prime}_{B}(u)$ where $\Gamma^{\prime}_{B}(u)=\frac{1}{P}\sum_{p=1}^{P}\exp{(ju\zeta^{\prime}_{p})}$. Now, since $\zeta_{p}\sim\mathcal{P}_{p}(\cdot)$, the random variable $\zeta^{\prime}_{p}$ has an even distribution because $\mathcal{P}_{p}(\cdot)$ is symmetric about $(P_{max}-1)/2$. Consequently, $\Gamma^{\prime}_{B}(u)$ is asymptotically jointly Gaussian distributed [57, Sec. II] as
[…]
@@
[…]
@@
Consider a MIMO radar with random SLA as described in Section II (of the main paper) with $\alpha_{1:N_{T}}$ and $\beta_{1:N_{R}}$ drawn i.i.d. from even distributions $\mathcal{P}_{\alpha}$ and $\mathcal{P}_{\beta}$, respectively. If $\mathcal{P}_{\alpha}$, $\mathcal{P}_{\beta}$ and the uniform grid $\phi_{1:G_{\theta}}$ satisfy (27), then for $0<\vartheta<1$ and $i=2,3,\ldots,G_{\theta}$,
@@
The lemma is obtained trivially by generalizing the results of [32, Theorem 1] to the MIMO-FMCW radar array. ∎
@@ Appendix C Proof of Theorem 2
According to [53, Theorem 2.7], any $K$-sparse signal can be recovered if the measurement matrix $\mathbf{D}$ has RIP constant $\delta_{2K}<2/(3+\sqrt{7/4})\doteq\Lambda$. Hence, we need to bound $\delta_{2K}$ with a probability higher than $1-\epsilon$, or equivalently, $\mathbb{P}(\delta_{2K}>\Lambda)<\epsilon$ where $0<\epsilon<1$. However, RIP constant $\delta_{2K}$ and coherence $\mu$ of a matrix satisfy $\delta_{2K}\leq(2K-1)\mu$ [53]. Hence, $\mathbb{P}(\delta_{2K}>\Lambda)\leq\mathbb{P}(\mu>\Lambda/(2K-1))$.
@@
$\times(1-2\vartheta_{0}\sqrt{N_{T}N_{R}}\;\mathcal{K}_{1}(2\vartheta_{0}\sqrt{N_{T}N_{R}}))^{G_{\theta}-1},$ (36)
@@
which can be solved using Lambert W function[60] as in the proof of [32, Theorem 2] to obtain (31). The constants $\kappa_{1}$ and $\kappa_{2}$ in (30)-(31) are obtained by substituting $\Lambda=2/(3+\sqrt{7/4})$.
For the case when $\alpha_{n}=\beta_{n}$, we have $\mathbb{P}(\delta_{2K}>\Lambda)<1-(1-e^{-\vartheta_{0}^{2}P})^{G_{D}-1}(1-e^{-N_{T}\vartheta_{0}})^{G_{\theta}-1}$ using (29) with $\vartheta_{0}=\Lambda/(2K-1)$. Again, using $(1-x)^{n}\approx 1-nx$ and $G_{D},G_{\theta}\gg 1$ and following similar steps as in the independent $\alpha_{1:N_{T}}$ and $\beta_{1:N_{R}}$ case, we obtain $G_{D}e^{-\vartheta_{0}^{2}P}+G_{\theta}e^{-N_{T}\vartheta_{0}}\leq\epsilon$, which is satisfied if $G_{D}e^{-\vartheta_{0}^{2}P}\leq\epsilon_{1}$ and $G_{\theta}e^{-N_{T}\vartheta_{0}}\leq\epsilon_{2}$ with $\epsilon_{1}+\epsilon_{2}=\epsilon$. Rearranging these inequalities, we obtain (30) and (32), respectively. Finally, the claim of the theorem then follows from [53, Theorem 2.7] considering the exact $K$-sparse signal case.
@@ Appendix D Proof of Theorem 3
If part: Using $\mathbf{D}_{N_{T}N_{R}(i-1)+j,:}=\mathbf{B}_{i,:}\otimes\mathbf{C}_{j,:}$, we obtain $\mathbb{E}[\mathbf{D}_{l,:}^{H}\mathbf{D}_{l,:}]=\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]\otimes\mathbb{E}[\mathbf{C}_{j,:}^{H}\mathbf{C}_{j,:}]$ for some $1\leq i\leq P$ and $1\leq j\leq N_{T}N_{R}$ because $\zeta_{1:P}$ is independent of $\alpha_{1:N_{T}}$ and $\beta_{1:N_{R}}$. We first consider $\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]$. Here, $\zeta_{1:P}$ are identically distributed such that $\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]$ does not depend on row index $i$ and $\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]=\frac{1}{P}\sum_{i=1}^{P}\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]$. But, by simple comparison of matrix elements, we can show that $\sum_{i=1}^{P}\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]=\mathbf{Q}_{B}$. Hence, $\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]=\widetilde{\mathbf{Q}}_{B}$ where $\widetilde{\mathbf{Q}}_{B}=(1/P)\mathbb{E}[\mathbf{Q}_{B}]$. As a consequence of Lemma 1, we only consider the first row of $\widetilde{\mathbf{Q}}_{B}$. Note that $(1,i)$-th element of $\mathbf{Q}_{B}$ is $\mathbf{B}_{:,1}^{H}\mathbf{B}_{:,i}=P\Gamma_{B}(u^{D}_{1,i})$ using (23). Hence, for $1\leq i\leq G_{D}$,
@@
where the second-last equality follows because $\zeta_{p}$ are i.i.d. for all $1\leq p\leq P$. Finally, if $\Psi_{p}(u^{D}_{1,i})=0$ for $i=2,3,\ldots,G_{D}$ according to (33), then $\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]=\widetilde{\mathbf{Q}}_{B}=\mathbf{I}_{G_{D}}$. Note that $\widetilde{\mathbf{Q}}_{B}]_{1,1}=1$ because all main diagonal elements of $\mathbf{Q}_{B}$ is $P$ from Lemma 1. Further, following similar steps as in proof of [32, Theorem 3], we can trivially show that $\mathbb{E}[\mathbf{C}_{j,:}^{H}\mathbf{C}_{j,:}]=\mathbf{I}_{G_{\theta}}$ if $\Psi_{\xi}(u^{\theta}_{1,j})=0$ for $j=2,3,\ldots,G_{\theta}$ as in (33). Finally, $\mathbb{E}[\mathbf{D}_{l,:}^{H}\mathbf{D}_{l,:}]=\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]\otimes\mathbb{E}[\mathbf{C}_{j,:}^{H}\mathbf{C}_{j,:}]=\mathbf{I}_{G_{D}}\otimes\mathbf{I}_{G_{\theta}}=\mathbf{I}_{G_{D}G_{\theta}}$, i.e., if (33) holds, matrix $\mathbf{D}$ satisfies isotropy property.
Only if part: If $\Psi_{p}(u^{D}_{1,i})\neq 0$ for some $i=2,3,\ldots,G_{D}$, then $\widetilde{\mathbf{Q}}_{B}]_{1,i}\neq 0$. Hence, $\mathbb{E}[\mathbf{B}_{i,:}^{H}\mathbf{B}_{i,:}]\neq\mathbf{I}$ which implies $\mathbb{E}[\mathbf{D}_{l,:}^{H}\mathbf{D}_{l,:}]\neq\mathbf{I}$, i.e., matrix $\mathbf{D}$ does not satisfy isotropy property. Similarly, matrix $\mathbf{D}$ also does not satisfy isotropy property if $\Psi_{\xi}(u^{\theta}_{1,j})\neq 0$ for some $j=2,3,\ldots,G_{\theta}$. Therefore, (33) is a necessary condition.
@@ Appendix E Proof of Proposition 5
Note that $\alpha_{1:N_{T}},\beta_{1:N_{R}}\in[-1/2,1/2]$ when $A_{T}=A_{R}=A/2$ and the characteristic function of uniform distributions $\mathcal{P}_{\alpha}$ and $\mathcal{P}_{\beta}$ are $\Psi_{\alpha}(u)=\Psi_{\beta}(u)=\frac{\sin(u/2)}{u/2}$. From (27), we need $\Psi_{\alpha}(u^{\theta}_{i,j})=\Psi_{\alpha}(2u^{\theta}_{i,j})=0$ and $\Psi_{\beta}(u^{\theta}_{i,j})=\Psi_{\beta}(2u^{\theta}_{i,j})=0$ for $i\neq j$ which can be satisfied if $u^{\theta}_{i,j}/2=k\pi$ for some integer $k$. But, by definition, $\frac{u^{\theta}_{i,j}}{2}=\pi\frac{A}{2\lambda}(\sin{\phi_{j}}-\sin{\phi_{i}})$. Hence, condition (27) is satisfied if $\phi_{1:G_{\theta}}$ is a uniformly spaced grid of spacing $2\lambda/A$ in the $\sin{\theta}$ domain.